OpenStudy (anonymous):

Show that the curve with parametric equations x=sint, y=cost, z=(sint)^2 is the curve of intersection of the surfaces z=x^2 and x^2 + y^2 =1. Use this fact to help sketch the curve

OpenStudy (anonymous):

no idea where to start, can someone guide me?

OpenStudy (nikvist):

use trigonometric identity \[\cos^2{\alpha}+\sin^2{\alpha}=1\]